### v1

# Upper bound on the multiplicity of rational and Du Bois singularities

Abstract.
This paper resolves a question of Huneke and Watanabe by proving a sharp upper bound for the multiplicity of Du Bois singularities: at a point of a $d$-dimensional variety with Du Bois singularities and embedding dimension $e$, the multiplicity is at most $\binom{e}{d}$. Additionally, the result recovers the previously known upper bound for the multiplicity of rational singularities.
2020 Mathematics Subject Classification:
13H15, 14B05
1. Introduction
[…]
The celebrated result of Kollár-Kovács [KK10] states that any union of log canonical centers has Du Bois singularities, and Kawamata’s subadjunction formula [Kawamata98] implies that a minimal log canonical center has rational singularities. Subsequently, Helmke’s multiplicity bound has been expected to be generalized to varieties with Du Bois or rational singularities. For a variety $X$ of dimension $d$ at a point $x\in X$ with embedding dimension $e=\dim m_{x}/m_{x}^{2}$, using commutative algebra methods, Huneke-Watanabe [HW15]*Theorem 3.1 proved the upper bound $\mathrm{mult}_{x}(X)\leq\binom{e-1}{d-1}$ when $X$ has rational singularities at $x\in X$, and asked if the upper bound $\mathrm{mult}_{x}(X)\leq\binom{e}{d}$ is true when $X$ has Du Bois singularities at $x\in X$ (see [HW15]*Remark 3.4). Our main theorem gives an affirmative answer to this question and, by the same methods, also recovers the bound for rational singularities.
Theorem 1.1.
Let $x\in X$ be a point in a variety with Du Bois (resp. rational) singularities. Denote $e:=\dim m_{x}/m_{x}^{2}$ and $d:=\dim X$ at $x$. Then we have
$\mathrm{mult}_{x}X\leq\binom{e}{d}\quad\left(\mathrm{resp.}\binom{e-1}{d-1}\right)$.
In positive characteristic, Huneke–Watanabe [HW15] proved the analogous bounds $\binom{e-1}{d-1}$ for F-rational singularities and $\binom{e}{d}$ for F-pure singularities. Note, however, that the bound for F-pure singularities does not imply that for Du Bois singularities in characteristic zero; see [Schwede09] for precise connections between the two settings. Shibata [Shibata17] proved the upper bound for a Cohen-Macaulay normal variety with Du Bois singularities, mainly following the proof of [HW15].
The proof of Theorem 1.1 is based on the multiplicity bound in $1$-dimensional case, which appears in Helmke’s work, with an argument arising from Koszul complexes in a resolution of singularities. Throughout this paper, a variety is a reduced separated scheme of finite type over $\mathbb{C}$. The embedding dimension of a variety $X$ at $x$ is the dimension of the cotangent space $m_{x}/m_{x}^{2}$ at $x$, where $m_{x}$ is the maximal ideal sheaf of $x\in X$.
Based on Theorem 1.1, we propose the following multiplicity bound in terms of minimal log discrepancies.
Conjecture 1.2.
Let $x\in X$ be a point on a variety with log canonical singularities. Denote $e:=\dim m_{x}/m_{x}^{2}$ and $d:=\dim X$ at $x$. If $\mathrm{mld}_{x}(X)>k-1$ for some integer $0\leq k\leq d$, then
$\mathrm{mult}_{x}X\leq\binom{e-k}{d-k}$.
Moreover, equality holds when k=d1k=d-1.
Here, $\mathrm{mld}_{x}(X)$ denotes the minimal log discrepancy of $X$ at $x$, i.e. the infimum of the log discrepancies $a(E;X)$ over all divisors $E$ with center $x$. Theorem 1.1 confirms the case $k=0$, and a minor refinement of its proof yields the case $k=1$. The equality statement is known in low dimensions: for surfaces with klt singularities by Artin [Artin66]*Corollary 6, and for threefolds with terminal singularities by Kakimi [Kakimi]*Theorem 2.1 and by Takagi–Watanabe [TW04]*Proposition 3.10.
2. Preliminaries
2.1. Multiplicities of points on curves
Let $X$ be a separated scheme of finite type over $\mathbb{C}$ and $x\in X$ be a closed point. For the multiplicity $\mathrm{mult}_{x}X$, it suffices to consider when $X$ has no embedded points through the following procedure. By the Noetherian property, there exists the largest ideal sheaf $\mathcal{I}\subset\mathcal{O}_{X}$ which is supported on a proper subvariety of $X$ not containing an irreducible component of $X$. Then the subscheme $X_{0}$ defined by $\mathcal{O}_{X}/\mathcal{I}$ satisfies $S_{1}$, or equivalently, there is no embedded point. It follows easily from the definition of the multiplicity that $\mathrm{mult}_{x}X=\mathrm{mult}_{x}X_{0}$.
Let $C$ be a curve (a separated scheme of finite type over $\mathbb{C}$ with dimension $1$) that satisfies $S_{1}$. Note that the blow-up $\mathrm{Bl}_{x}C$ of a closed point $x\in C$ satisfies $S_{1}$. Indeed, the blow-up $\mathrm{Bl}_{x}C$ is the scheme-theoretic closure of $C\setminus\left\{x\right\}$ in $\mathrm{Bl}_{x}C$, so that $\mathrm{Bl}_{x}C$ has no embedded point if $C\setminus\left\{x\right\}$ has no embedded point.
Lemma 2.1 ([Helmke97]*Lemma 4.2).
Let $C$ be a curve with no embedded point. Let $\mu:\mathrm{Bl}_{x}C\to C$ be the blow-up of a closed point $x\in C$ and $E_{x}:=m_{x}\mathcal{O}_{\mathrm{Bl}_{x}C}$ be the exceptional divisor. Assume that for some $k\geq 0$, we have a natural inclusion
$\mu_{*}\mathcal{O}_{\mathrm{Bl}_{x}C}(-kE_{x})\subset\mathcal{O}_{C}$
in $\mathcal{R}_{C}:=\oplus_{\eta}\mathcal{O}_{C,\eta}$, the sheaf of rational functions of $C$ where $\eta$ ranges over the generic points of $C$. Then
$\mathrm{mult}_{x}C\leq\binom{e+k-1}{k}$
where $e=\dim m_{x}/m_{x}^{2}$ is the embedding dimension of $C$ at $x$.
Note that $\mu:\mathrm{Bl}_{x}C\to C$ is an isomorphism over $C\setminus\left\{x\right\}$, which implies that both $\mu_{*}\mathcal{O}_{\mathrm{Bl}_{x}C}(-kE_{x})$ and $\mathcal{O}_{C}$ are subsheaves of $\mathcal{R}_{C}$. We include the proof of Helmke [Helmke97] for the reader’s convenience.
Proof.
Let $V=m_{x}/m_{x}^{2}$ be the cotangent space of $C$ at $x$. By choosing $e$ elements of $m_{x}$ whose residue classes form a basis of the vector space $V$, we obtain a surjection of rings
$\Phi:\bigoplus^{k}_{i=0}Sym^{i}V\to\mathcal{O}_{C}/m_{x}^{k+1}.$
Consider the following Cartesian square:
[…]
Since $\Phi$ is surjective, the first column is surjective. The elements in $m_{x}^{k+1}$ are sections of
$\mu_{*}\mathcal{O}_{\mathrm{Bl}_{x}C}(-(k+1)E_{x}),$
and thus we have a surjection
$\phi:I\to\mu_{*}\left(\mathcal{O}_{\mathrm{Bl}_{x}C}(-kE_{x})/\mathcal{O}_{\mathrm{Bl}_{x}C}(-(k+1)E_{x})\right)\cong\mu_{*}\mathcal{O}_{E_{x}}(-kE_{x}).$
Note that for any nonzero element $t\in V$, $tI$ is contained in the kernel of $\phi$. Therefore
$\mathrm{mult}_{x}C=h^{0}(E_{x},\mathcal{O}_{E_{x}}(-kE_{x}))\leq\dim I-\dim tI=\dim\left(\ker(t:I\to tI)\right)$
and $\ker(t:I\to tI)=Sym^{k}V$, from which we conclude that $\mathrm{mult}_{x}C\leq\binom{e+k-1}{k}$. ∎
2.2. Koszul complexes associated to sets of Cartier divisors
Let $\mathcal{E}$ be a vector bundle of rank $r$ on $X$, a separated scheme of finite type over $\mathbb{C}$. Given a morphism $\phi:\mathcal{E}\to\mathcal{O}_{X}$, there exists a Koszul complex $K_{\bullet}(\phi)$:
$0\to\wedge^{r}\mathcal{E}\to\dots\to\mathcal{E}\to\mathcal{O}_{X}\to 0.$
Locally, let $e_{1},\dots,e_{r}$ be a generator of $\mathcal{E}$, then each map $d:\wedge^{j}\mathcal{E}\to\wedge^{j-1}\mathcal{E}$ for $1\leq j\leq r$ is described by
$d(e_{i_{1}}\wedge\cdots\wedge e_{i_{j}})=\sum_{k=1}^{j}(-1)^{k+1}\phi(e_{i_{k}})e_{i_{1}}\wedge\cdots\wedge\widehat{e_{i_{k}}}\wedge\cdots\wedge e_{i_{j}}$
where $e_{i_{1}}\wedge\cdots\wedge e_{i_{j}}$ is a generator of $\wedge^{j}\mathcal{E}$.
We consider the case when $\mathcal{E}=\oplus_{i=1}^{r}\mathcal{O}_{X}(-D_{i})$ such that $D_{i}$ is an effective Cartier divisor, and $\phi:\mathcal{E}\to\mathcal{O}_{X}$ is the sum of natural inclusions $\mathcal{O}_{X}(-D_{i})\to\mathcal{O}_{X}$. Note that
$\wedge^{j}\mathcal{E}\cong\bigoplus_{1\leq i_{1}<\dots<i_{j}\leq r}\mathcal{O}_{X}(-D_{i_{1}}-\cdots-D_{i_{j}}).$
Consider the complex
$\cdots\to\bigoplus_{1\leq i_{1}<\dots<i_{j}\leq r}\mathcal{O}_{X}(-D_{i_{1}}-\cdots-D_{i_{j}})\xrightarrow{d}\bigoplus_{1\leq i_{1}<\dots<i_{j-1}\leq r}\mathcal{O}_{X}(-D_{i_{1}}-\cdots-D_{i_{j-1}})\to\cdots,$
whose differential $d$ on each component $\mathcal{O}_{X}(-D_{i_{1}}-\cdots-D_{i_{j}})$ is given by the alternating sum $\sum_{k=1}^{j}(-1)^{k+1}\iota_{k}$ of natural inclusions
$\iota_{k}:\mathcal{O}_{X}(-D_{i_{1}}-\cdots-D_{i_{j}})\to\mathcal{O}_{X}(-D_{i_{1}}-\cdots-\widehat{D_{i_{k}}}-\cdots-D_{i_{j}}).$
Recall that $\mathcal{O}_{X}(-D_{i_{1}}-\cdots-D_{i_{j}})$ is the sub-$\mathcal{O}_{X}$-module of the sheaf of total quotient rings of $\mathcal{O}_{X}$ and $\iota_{k}$ is induced by this inclusion. Then this complex is the Koszul complex $K_{\bullet}(\phi)$.
Indeed, let $f_{1},\dots,f_{r}$ be the local generators of $D_{1},\dots,D_{r}$. Then $\mathcal{O}_{X}(-D_{i_{1}}-\cdots-D_{i_{j}})$ represents an ideal sheaf generated by $f_{i_{1}}\cdots f_{i_{j}}$. Treating the generator $f_{i_{1}}\cdots f_{i_{j}}$ as $e_{i_{1}}\wedge\cdots\wedge e_{i_{j}}$, we recover the local description of the Koszul complex $K_{\bullet}(\phi)$.
In the rest of the paper, we make use of this Koszul complex obtained from a set of Cartier divisors $D_{1},\dots,D_{r}$. We follow the convention that for $j\geq 0$, the $j-$th degree component of $K_{\bullet}(\phi)$ is
$K_{-j}(\phi):=\bigoplus_{1\leq i_{1}<\dots<i_{j}\leq r}\mathcal{O}_{X}(-D_{i_{1}}-\cdots-D_{i_{j}}).$
3. Proof of Theorem 1.1 for rational singularities
Let $x\in X$ be a point in a variety with rational singularities, $f:\mathrm{Bl}_{x}X\to X$ be the blow-up at $x$, and $g:\widetilde{X}\to\mathrm{Bl}_{x}X$ be a resolution of singularities. Denote by $\mu:=f\circ g:\widetilde{X}\to X$, and let $\widetilde{E}_{x}:=g^{*}E_{x}$ be the pullback of the exceptional divisor of the blow-up $f$.
For a general hyperplane section $H$ of $X$ through $x$, its strict transform $\widetilde{H}$ in $\widetilde{X}$ is a base point free Cartier divisor such that $\mu^{*}H=\widetilde{H}+\widetilde{E}_{x}$. Since $X$ has rational singularities, we have $R\mu_{*}\mathcal{O}_{\widetilde{X}}=\mathcal{O}_{X}$, which induces the natural identity
(3.1) $R\mu_{*}\mathcal{O}_{\widetilde{X}}(-\widetilde{H}-\widetilde{E}_{x})=\mathcal{O}_{X}(-H)$
from the projection formula.
Take general hyperplane sections $H_{1},\dots,H_{d-1}$ through $x\in X$. Denote $C:=H_{1}\cap\cdots\cap H_{d-1}$. Then we have $\mathrm{mult}_{x}X=\mathrm{mult}_{x}C$. Consider the Koszul complex $K_{\bullet}(\phi)$ obtained from the morphism $\phi:\oplus_{i=1}^{d-1}\mathcal{O}_{\widetilde{X}}(-\widetilde{H}_{i}-\widetilde{E}_{x})\to\mathcal{O}_{\widetilde{X}}$ and the Koszul complex $K_{\bullet}(\psi)$ obtained from the morphism $\psi:\oplus_{i=1}^{d-1}\mathcal{O}_{\widetilde{X}}(-\widetilde{H}_{i})\to\mathcal{O}_{\widetilde{X}}$. Since $\widetilde{H}$ is base point free, $\widetilde{H}_{1},\dots,\widetilde{H}_{d-1}$ is a regular sequence and their intersection $\widetilde{C}:=\widetilde{H}_{1}\cap\cdots\cap\widetilde{H}_{d-1}$ is a smooth curve which maps surjectively to $C$. Therefore, we have $K_{\bullet}(\psi)\cong\mathcal{O}_{\widetilde{C}}$.
Consider the twisted complex $K_{\bullet}(\psi)\otimes\mathcal{O}_{\widetilde{X}}(-(d-1)\widetilde{E}_{x})$. Then for $0\leq j\leq d-1$, we have
$K_{-j}(\psi)\otimes\mathcal{O}_{\widetilde{X}}(-(d-1)\widetilde{E}_{x})=\bigoplus_{1\leq i_{1}<\dots<i_{j}\leq d-1}\mathcal{O}_{\widetilde{X}}(-\widetilde{H}_{i_{1}}-\cdots-\widetilde{H}_{i_{j}}-(d-1)\widetilde{E}_{x}).$
Then we have a natural inclusion
$K_{-j}(\psi)\otimes\mathcal{O}_{\widetilde{X}}(-(d-1)\widetilde{E}_{x})\to K_{-j}(\phi)$
induced by the natural inclusion
$\mathcal{O}_{\widetilde{X}}(-\widetilde{H}_{i_{1}}-\cdots-\widetilde{H}_{i_{j}}-(d-1)\widetilde{E}_{x})\to\mathcal{O}_{\widetilde{X}}(-\widetilde{H}_{i_{1}}-\cdots-\widetilde{H}_{i_{j}}-j\widetilde{E}_{x})$
of subsheaves of the sheaf of the total quotient ring of $\mathcal{O}_{\widetilde{X}}$ for each summand. From the description of Koszul complexes in Section 2.2, this induces a morphism of complexes
(3.2) $K_{\bullet}(\psi)\otimes\mathcal{O}_{\widetilde{X}}(-(d-1)\widetilde{E}_{x})\to K_{\bullet}(\phi).$
Recall that $K_{\bullet}(\psi)\otimes\mathcal{O}_{\widetilde{X}}(-(d-1)\widetilde{E}_{x})\cong\mathcal{O}_{\widetilde{C}}(-(d-1)\widetilde{E}_{x})$. Additionally, we have the natural isomorphism of complexes
$R\mu_{*}(K_{\bullet}(\phi))\cong K_{\bullet}(\oplus_{i=1}^{d-1}\mathcal{O}_{X}(-H_{i})\to\mathcal{O}_{X})$
from (3.1). Therefore, $R^{0}\mu_{*}$ of (3.2) induces a morphism
$\mu_{*}\mathcal{O}_{\widetilde{C}}(-(d-1)\widetilde{E}_{x})\to\mathcal{O}_{C}.$
Note that $\mu:\widetilde{C}\to C$ factors through the blow-up $\mathrm{Bl}_{x}C$, which induces a natural morphism
$\mathcal{O}_{\mathrm{Bl}_{x}C}(-(d-1)E_{x})\to g_{*}\mathcal{O}_{\widetilde{C}}(-(d-1)\widetilde{E}_{x}).$
Consequently, we obtain a morphism
(3.3) $f_{*}\mathcal{O}_{\mathrm{Bl}_{x}C}(-(d-1)E_{x})\to\mathcal{O}_{C}.$
Here, $C$ has no embedded point since it is cut out by a regular sequence of hyperplane sections $H_{1},\dots,H_{d-1}$ on a Cohen-Macaulay variety $X$.
Going back to (3.2), we have a commuting diagram of morphisms of complexes
${{\mathcal{O}_{\widetilde{X}}(-(d-1)\widetilde{E}_{x})}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{O}_{\widetilde{X}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{K_{\bullet}(\psi)\otimes\mathcal{O}_{\widetilde{X}}(-(d-1)\widetilde{E}_{x})}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{K_{\bullet}(\phi)}}$
where the top row is obtained by taking $\bullet=0$ of the bottom row. Suppose a local section $s$ of $f_{*}\mathcal{O}_{\mathrm{Bl}_{x}C}(-(d-1)E_{x})$ maps to a local section $s^{\prime}$ of $\mathcal{O}_{C}$ via (3.3). Then $s^{\prime}$ lifts to a section of $\mathcal{O}_{\widetilde{X}}$ which can be considered as a rational section of $\mathcal{O}_{\widetilde{X}}(-(d-1)\widetilde{E}_{x})$. This rational section restricts back to the section $s$ of $\mathcal{O}_{\mathrm{Bl}_{x}C}(-(d-1)E_{x})$, because (3.2) is a natural isomorphism over $X\setminus x$. Therefore, (3.3) is a natural inclusion of subsheaves of the sheaf of rational functions of $C$.
The embedding dimension of $C$ at $x$ is ed+1e-d+1. Therefore, by Lemma 2.1, we obtain the multiplicity bound $\binom{e-1}{d-1}$ as desired.
4. Proof of Theorem 1.1 for Du Bois singularities
Let $x\in X$ be a point in a variety with Du Bois singularities. In the neighborhood of $x$, consider a closed embedding $X\subset W$ into a smooth variety $W$. Let $f:\mathrm{Bl}_{x}W\to W$ be the blow-up of $W$ at $x$, and $g:\widetilde{W}\to\mathrm{Bl}_{x}W$ be a log resolution of singularities such that the set-theoretic pullback $(f\circ g)^{-1}(X)=E+F$ is a reduced simple normal crossing divisor with $E=\mu^{-1}(x)$. Denote by $\mu:=f\circ g:\widetilde{W}\to W$, and let $\widetilde{E}_{x}:=g^{*}E_{x}$ be the pullback of the exceptional divisor of the blow-up $f$. Then $E$ is a reduced divisor of $\widetilde{E}_{x}$.
As in the proof for rational singularities, choose a general hyperplane section $H$ of $W$ through $x$. Then $\mu^{*}H=\widetilde{H}+\widetilde{E}_{x}$, where $\widetilde{H}$ is a base point free Cartier divisor. Since $W$ is smooth, we have an identity
(4.1) $R\mu_{*}\mathcal{O}_{\widetilde{W}}(-\widetilde{H}-\widetilde{E}_{x})=\mathcal{O}_{W}(-H).$
Additionally, $\mu$ is an isomorphism over the complement of $X$. Thus, we have a natural quasi-isomorphism of Du Bois complexes
$R\mu_{*}\underline{\Omega}_{\widetilde{W},E\cup F}^{0}=\underline{\Omega}_{W,X}^{0}.$
See, for example, Du Bois [DuBois81]*Proposition 4.11. Here, $\underline{\Omega}^{0}_{W,X}$ is the mapping cone of the natural morphism $\rho:\underline{\Omega}^{0}_{W}\to\underline{\Omega}^{0}_{X}$ of Du Bois complexes shifted by [1][-1]. In particular, if both $W$ and $X$ are Du Bois, then $\underline{\Omega}^{0}_{W,X}=\mathcal{I}_{W,X}$ where $\mathcal{I}_{W,X}$ is the ideal sheaf of $X\subset W$. This implies the identity
(4.2) $R\mu_{*}\mathcal{O}_{\widetilde{W}}(-F-E)=\mathcal{I}_{W,X}$
(see also [Park23]*Corollary 1.10).
Take general hyperplane sections $H_{1},\dots,H_{d-1}$ through $x\in W$, and denote by $C:=X\cap H_{1}\cap\cdots\cap H_{d-1}$. Then we have $\mathrm{mult}_{x}X=\mathrm{mult}_{x}C$. Consider the Koszul complex $K_{\bullet}(\phi)$ obtained from the morphism
$\phi:\mathcal{O}_{\widetilde{W}}(-F-E)\oplus\bigoplus_{i=1}^{d-1}\mathcal{O}_{\widetilde{W}}(-\widetilde{H}_{i}-\widetilde{E}_{x})\to\mathcal{O}_{\widetilde{W}}$
and the Koszul complex $K_{\bullet}(\psi)$ obtained from the morphism
$\psi:\mathcal{O}_{\widetilde{W}}(-F)\oplus\bigoplus_{i=1}^{d-1}\mathcal{O}_{\widetilde{W}}(-\widetilde{H}_{i})\to\mathcal{O}_{\widetilde{W}}$.
Note that $F,\widetilde{H}_{1},\dots,\widetilde{H}_{d-1}$ is a regular sequence and their intersection $V:=F\cap\widetilde{H}_{1}\cap\cdots\cap\widetilde{H}_{d-1}$ is a reduced simple normal crossing divisor in a smooth variety $\widetilde{H}_{1}\cap\cdots\cap\widetilde{H}_{d-1}$. In particular, we have $K_{\bullet}(\psi)\cong\mathcal{O}_{V}$ and it is easy to see that $V$ is surjectively mapped to $C$.
Consequently, we have a morphism of complexes
(4.3) $K_{\bullet}(\psi)\otimes\mathcal{O}_{\widetilde{W}}(-d\widetilde{E}_{x})\to K_{\bullet}(\phi),$
induced by the natural inclusions
$\mathcal{O}_{\widetilde{W}}(-\widetilde{H}_{i_{1}}-\cdots-\widetilde{H}_{i_{j}}-d\widetilde{E}_{x})\to\mathcal{O}_{\widetilde{W}}(-\widetilde{H}_{i_{1}}-\cdots-\widetilde{H}_{i_{j}}-j\widetilde{E}_{x})$
and
$\mathcal{O}_{\widetilde{W}}(-F-\widetilde{H}_{i_{1}}-\cdots-\widetilde{H}_{i_{j}}-d\widetilde{E}_{x})\to\mathcal{O}_{\widetilde{W}}(-F-\widetilde{H}_{i_{1}}-\cdots-\widetilde{H}_{i_{j}}-E-j\widetilde{E}_{x})$
of subsheaves of the sheaf of the total quotient ring of $\mathcal{O}_{\widetilde{W}}$ for each $0\leq j\leq d$.
From (4.1) and (4.2), we have $R^{0}\mu_{*}(K_{\bullet}(\phi))=\mathcal{O}_{C}$. Recall that $K_{\bullet}(\psi)\otimes\mathcal{O}_{\widetilde{W}}(-d\widetilde{E}_{x})\cong\mathcal{O}_{V}(-d\widetilde{E}_{x})$. Therefore, $R^{0}\mu_{*}$ of (4.3) induces a morphism
$\mu_{*}\mathcal{O}_{V}(-d\widetilde{E}_{x})\to\mathcal{O}_{C}.$
Since $\mu:V\to C$ factors through the blow-up $\mathrm{Bl}_{x}C$, we have a morphism
$f_{*}\mathcal{O}_{\mathrm{Bl}_{x}C}(-dE_{x})\to\mathcal{O}_{C}.$
Note that by the general choice of hyperplane sections, $C$ is reduced away from $x$. Let $C_{red}$ be the reduced scheme of $C$. Then we deduce a morphism
(4.4) $f_{*}\mathcal{O}_{\mathrm{Bl}_{x}C}(-dE_{x})\to\mathcal{O}_{C_{red}}.$
Consider a commuting diagram of morphisms of complexes
${{\mathcal{O}_{\widetilde{W}}(-d\widetilde{E}_{x})}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathcal{O}_{\widetilde{W}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{K_{\bullet}(\psi)\otimes\mathcal{O}_{\widetilde{W}}(-d\widetilde{E}_{x})}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{K_{\bullet}(\phi)}}$
where the top row is obtained by taking $\bullet=0$ of the bottom row. Suppose a local section $s$ of $f_{*}\mathcal{O}_{\mathrm{Bl}_{x}C}(-dE_{x})$ maps to a local section $s^{\prime}$ of $\mathcal{O}_{C_{red}}$ via (4.4). Then $s^{\prime}$ lifts to a section of $\mathcal{O}_{\widetilde{W}}$ which can be considered as a rational section of $\mathcal{O}_{\widetilde{W}}(-d\widetilde{E}_{x})$. This rational section restricts to a rational section $s_{V}$ of $\mathcal{O}_{V}(-d\widetilde{E}_{x})$. Because (4.3) is a natural isomorphism over $W\setminus x$, the rational section $s_{V}$ is equal to the section $s$ pulled back to $V$ over the complement of $x$. Therefore, $s$ and $s^{\prime}$ are the same over the complement of $x$, which indicates that (4.4) is a natural inclusion of subsheaves of the sheaf of rational functions of $C_{red}$.
The embedding dimension of $C$ at $x$ is ed+1e-d+1. Therefore, by Lemma 2.1, we obtain the multiplicity bound $\binom{e}{d}$ as desired.
5. Boundary examples
In this section, we provide an explicit example of a Du Bois singularity $x\in X$ of dimension $d$ and embedding dimension $e$ with $\mathrm{mult}_{x}X=\binom{e}{d}$. Let $\mathbb{A}^{e}$ be the affine space with coordinates $x_{1},\dots,x_{e}$. Let $X\subset\mathbb{A}^{e}$ be the union of $d$-dimensional planes spanned by $d$ coordinate axes. Hence, $X$ is a union of $\binom{e}{d}$ planes of dimension $d$. It is well known and easy to see using inductive arguments that $X$ has Du Bois singularities, and $\mathrm{mult}_{x}X=\binom{e}{d}$ as desired. See [HW15]*Section 4 for a boundary example for rational singularities.

### v2

@@
 The embedding dimension of a variety $X$ at $x$ is the dimension of the cotangent space $m_{x}/m_{x}^{2}$ at $x$, where $m_{x}$ is the maximal ideal sheaf of $x\in X$.
-Based on Theorem 1.1, we propose the following multiplicity bound in terms of minimal log discrepancies.
+Motivated by Theorem 1.1 and Shibata [Shibata17a]*Proposition 3.9, we propose the following multiplicity bound in terms of minimal log discrepancies.
 Conjecture 1.2.
@@
 the infimum of the log discrepancies $a(E;X)$ over all divisors $E$ with center $x$.
+Note that [Shibata17a] proves the statement when $X$ is klt and $\mathrm{mld}_{x}(X)$ is replaced by the log canonical threshold of the maximal ideal at $x\in X$.
 Theorem 1.1 confirms the case $k=0$, and a minor refinement of its proof yields the case $k=1$.
-The equality statement is known in low dimensions: for surfaces with klt singularities by Artin [Artin66]*Corollary 6, and for threefolds with terminal singularities by Kakimi [Kakimi]*Theorem 2.1 and by Takagi–Watanabe [TW04]*Proposition 3.10.
+The equality statement is known in low dimensions: for surfaces with klt singularities by Artin [Artin66]*Corollary 6, and for threefolds with terminal singularities by Kakimi [Kakimi]*Theorem 2.1 and by Takagi-Watanabe [TW04]*Proposition 3.10.
 2. Preliminaries